# Sensor-fusion based Prognostics Framework for Complex Engineering Systems Exhibiting Multiple Failure Modes

### v2

@@ 6.3 RUL Prediction
Each of the datasets consist of 80 test systems (40 from each failure mode). For each test system, we observe degradation up until the following life percentiles (in retrospect): 10%, 20%, 30%, 40%, 50%, 60%, 70%, 80%, and 90%. We smooth the training and the test signals using a bandwidth parameter of 0.5. Next, we perform MFPCA selecting the number using only the training systems that survived up to the observation time $t^{*}$ corresponding to each life percentile. For a given test signal, we calculate its MFPC scores. The active failure mode is diagnosed by being assigned to the most frequent failure mode present in its closest $0.1N_{t^{*}}$ neighbors. Once the active failure mode is identified, the MFPC scores of the test signals are updated and recomputed using the subset of training signals associated with the active failure mode. Next, we fit the weighted regression model and predict the RUL using the scores of the test signal. Prediction accurately is calculated retrospectively. We compute the “Estimated Life” as the current operating time plus the predicted RUL. A relative error is then computed using equation (21).
@@ 7 Case Study 2: NASA Turbofan Engine Dataset
To further validate our methodology, we evaluate its performance on one of the Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) datasets [44]. C-MAPSS is a simulation module that generates turbofan engine degradation datasets. Figure 4 displays a diagram of this turbofan engine. In this paper, we utilize C-MAPSS-3, which consists of both a training and a test set. Both sets contain 100 observed engine failures caused by either fan degradation or high-pressure compressor degradation. Furthermore, both sets contain data from 21 sensors, which are listed in Table 4. The signals for these sensors are displayed in Figure 5. These sensors monitor the systems from a good-as-new state to engine failure. However, sensors 1, 5, 6, 10, 16, 18, and 19 contain little to no information and are thus removed from the analysis. While the dataset provides the potential causes of failure, it does not provide the failure labels. Furthermore, the test set consists of systems whose degradation is only partially observed along with their corresponding failure times. Therefore, our goal is to use our methodology to predict the remaining life of each testing unit following the final cycle at which their degradation was observed.
@@ 21 W32 LPT coolant bleed
Benchmarking: To analyze our methodology, we compare its performance to that of [32] (Chehade et al. 2018), whose methodology assumes knowledge of the failure modes but does not perform sensor selection (outside of removing the blank sensors), and [33] (Li et al. 2023), whose methodology assumes knowledge of a fraction of the training set and does perform sensor selection. Figure 6 displays the relative prediction error averaged over all test systems with RUL at most 100, 80, 60, 40, and 20 cycles. From this figure, our methodology becomes more accurate as the RUL decreases. Furthermore, our methodology is more accurate than both when making longer predictions. For test units whose RUL is at most 20, our accuracy is between that of [32] and [33]. Only at RUL at most 40, does our methodology perform worse than both. To further analyze our methodology, we refer to Figure 7, which displays boxplots showing the distribution of the relative prediction error for test units with remaining life in the following intervals: $[0,25],[26,50],[51,75],[76,100],[101,125],[126,150]$. The results indicate that our methodology achieves a lower median relative error than [32] in all intervals while maintaining comparable variability. However, the variability for systems with RUL between 25 and 50 is higher for our methodology than for [32]. This indicates a targeted area for improvement of our methodology. However, the general trend is that sensor selection resulted in a more accurate RUL prediction than no sensor selection. Furthermore, we were capable of achieving more accurate RUL prediction than both methods from earlier in the systems’ lives despite not having information about the actual failure mode.
@@ 8 Conclusion
In this paper, we present a methodology for predicting the RUL of systems with multiple failure modes using multiple sensors. This methodology consists of two phases: an offline phase and an online phase. Assuming that failure modes are mutually exclusive, the offline phase consists of a framework for diagnosing the observed failures in the training set and selecting an optimal subset of sensors for monitoring. In this framework, we utilize CA-FPCA to capture the multiple failure modes exhibited by the sensor signals enabling the extraction of informative features. These features are used by our proposed EM algorithm to train the MGR-ASGL model which simultaneously diagnoses the training set failures and selects an optimal subset of informative sensors for each failure mode. In the online phase, we use MFPCA to fuse all the signals of the selected sensors into a parsimonious set of features to diagnose the failure mode for an actively degrading system. Then MFPCA is re-used to fuse the signals from informative sensors of the diagnosed failure mode for RUL prediction. RUL prediction is done using a time-varying weighted functional regression model.
To assess the performance of our methodology, we implemented it on two datasets – a) a simulated dataset and b) the NASA turbofan engine degradation dataset. For the simulated dataset, we tested the robustness of our methodology to various degrees of signal noise. We demonstrated that with a sufficiently high $\lambda$ value, the clustering matches the actual clusters at an accuracy of about 80%. Furthermore, the most significant sensors occupied 3 of the top 4 slots with regard to $l_{2}$-norm of regression coefficients. Finally, the methodology showed high accuracy with an average relative prediction error of no larger than 3, irrespective of the percentage of lifetime the prediction was performed.
In our real-world data case study, results show that our model makes early predictions more accurately than the two benchmarks while achieving comparable accuracy in later predictions. To achieve this, we identified 10 sensors to monitor the first failure mode while identifying 8 sensors to monitor the second failure mode. Our methodology removed one sensor completely. The region in which our methodology performed worst relative to the two baselines was when the RUL was between 25 and 50. A possibility for improvement would be to relax the assumption of a linear relationship between the $lnTTF$ and the sensor signals. However, embedding a variable selection technique within a nonlinear regression model is an open topic in the literature. Therefore, the extension of this work to a nonlinear setting is a topic for future work.

### v3

@@ 6.3 RUL Prediction
Each of the datasets consist of 80 test systems (40 from each failure mode). For each test system, we observe degradation up until the following life percentiles (in retrospect): 10%, 20%, 30%, 40%, 50%, 60%, 70%, 80%, and 90%. We smooth the training and the test signals using a bandwidth parameter of 0.5. Next, we perform MFPCA selecting the number using only the training systems that survived up to the observation time $t^{*}$ corresponding to each life percentile. For a given test signal, we calculate its MFPC scores. The active failure mode is diagnosed by being assigned to the most frequent failure mode present in its closest $0.1N_{t^{*}}$ neighbors. Once the active failure mode is identified, the MFPC scores of the test signals are updated and recomputed using the subset of training signals associated with the active failure mode. Next, we fit the weighted regression model and predict the RUL using the scores of the test signal. Prediction accurately is calculated retrospectively. We compute the “Estimated Life” as the current operating time plus the predicted RUL. A relative error is then computed using Eq (24).
@@ 7 Case Study 2: NASA Turbofan Engine Dataset
To further validate our methodology in a realistic operational context, we evaluate its performance on a benchmark dataset that closely mirrors subsystem degradation in deep-space habitats. Specifically, we apply our approach to the NASA’s Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) dataset [55]. C-MAPSS simulates degradation in the components of turbofan engines that share telemetry and fault characteristics with propulsion and environmental control systems of deep space habitats. C-MAPSS involves multivariate sensor monitoring from healthy operation to system failure, with no explicit failure mode labels and only partial degradation histories for test systems. Figure 5 displays a diagram of this turbofan engine.
In this paper, we utilize C-MAPSS-3, which consists of both a training and a test set. Both sets contain 100 observed engine failures caused by either fan degradation or high-pressure compressor degradation. Furthermore, both sets contain data from 21 sensors, which are listed in Table 5. The signals for these sensors are displayed in Figure 6. These sensors monitor the systems from a good-as-new state to engine failure. However, sensors 1, 5, 6, 10, 16, 18, and 19 contain little to no information and are thus removed from the analysis. While the dataset provides the potential causes of failure, it does not provide the failure labels. Furthermore, the test set consists of systems whose degradation is only partially observed along with their corresponding failure times. Therefore, our goal is to use our methodology to predict the remaining life of each testing unit following the final cycle at which their degradation was observed.
@@ 21 W32 LPT coolant bleed
7.1 Comparison to Baselines
To analyze our methodology, we compare its performance to that of [39] Chehade et al. (2018), whose methodology assumes knowledge of the failure modes but does not perform sensor selection (outside of removing the blank sensors), and [40] Li et al. (2023), whose methodology assumes knowledge of a fraction of the training set and does perform sensor selection. Figure 7 displays the relative prediction error averaged over all test systems with RUL at most 100, 80, 60, 40, and 20 cycles. From this figure, our methodology becomes more accurate as the RUL decreases. Furthermore, our methodology is more accurate than both when making longer predictions. For test units whose RUL is at most 20, our accuracy is between that of [39] and [40]. Only at RUL at most 40, does our methodology perform worse than both. To further analyze our methodology, we refer to Figure 8, which displays boxplots showing the distribution of the relative prediction error for test units with remaining life in the following intervals: $[0,25],[26,50],[51,75],[76,100],[101,125],[126,150]$. The results indicate that our methodology achieves a lower median relative error than [39] in all intervals while maintaining comparable variability. However, the variability for systems with RUL between 25 and 50 is higher for our methodology than for [39]. This indicates a targeted area for improvement of our methodology. However, the general trend is that sensor selection resulted in a more accurate RUL prediction than no sensor selection. Furthermore, we were capable of achieving more accurate RUL prediction than both methods from earlier in the systems’ lives despite not having information about the actual failure mode.
@@ 8 Conclusion
In this paper, we presented a prognostics methodology tailored to the unique challenges of deep-space habitats. Complex engineering systems such as deep space habitats, are monitored by a large number of sensors and are susceptible to multiple failure modes that are often unlabeled. Our approach addresses the need for autonomous and accurate remaining useful life prediction in environments where expert diagnosis and ground-based maintenance are not feasible.
The methodology consists of two phases. In the offline phase, we extract failure-mode-specific features using component-aligned functional principal component analysis. These features are used by a novel expectation maximization algorithm to train a regression model that simultaneously clusters systems by failure mode and selects the most informative sensors for each mode. In the online phase, we use multivariate functional principal component analysis to reduce the dimensionality of real-time sensor data, diagnose the active failure mode, and predict remaining useful life using a time-varying weighted functional regression model.
We validated our methodology using two case studies. The first is a simulated dataset that reflects key properties of deep-space habitats, including mode-specific degradation behavior, redundant sensors, and variable signal-to-noise ratios. Our method achieved strong performance in clustering failure modes, identifying informative sensors, and predicting remaining useful life with low relative error. The second case study applied our method to the C-MAPSS dataset, which we used to test real-world feasibility under partial degradation observation and unlabeled failure conditions. Our model outperformed baseline approaches in early-life prediction and produced interpretable results by selecting a minimal and relevant set of sensors.
This work is based on the assumption of a linear relationship between the log of time-to-failure (TTF) and the sensor signals. A possibility for improvement would be to relax this assumption. However, embedding a variable selection technique within a nonlinear regression model is an open topic in the literature. Future work will focus on extending the methodology to nonlinear models to improve prediction accuracy and capture complex degradation trends in autonomous mission scenarios.
@@ Acknowledgments
Appendix A Log Likelihood
Definition 1.
Definition 2.
$\begin{split}\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})&=\sum_{i=1}^{N}\sum_{k=1}^{K}\boldsymbol{Z}_{i}[k]\ln\pi_{k}\left(\frac{\rho_{k}}{\sqrt{2\pi}}\right)\times\\ &\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)\end{split}$ (26)
Lemma 1.
Proof.
where,
Since $\ln(.)$ is concave, from Jensen’s inequality we can write that,
and thus,